A local financial consulting firm employs 20 full-time staff. The estimated compensation per employee is $50838 for 2049 hours. All direct labour costs are charged to clients. Any other costs are included in a single indirect-cost pool, allocated according to labour-hours.

If the total cost of a job that will take 23 hours is $2447 using a job order cost system, what is the budgeted indirect costs for the coming year?

Select one:

a. $1876

b. $3343156

c. $3997143

d. $104167062

Data regarding Lager Ltd.’s production for the year is given below:

Per unit information:

Direct labour

$13

Direct material

9

Applied manufacturing overhead

8

Unit production (budget)

31184

Unit production (actual)

23971

Actual overhead costs incurred

$132674

Lager Ltd.’s manufacturing overhead for the year was:

Select one:

a. Under-applied by $59094

b. Under-applied by $116798

c. Over-applied by $116798

d. Over-applied by $59094
Answer #1

Please provide rating...

Answer 1
Total cost = 2447
Direct cost per hr =        24.81
=50838/2049
Direct cost of the job 571
=24.81*23
Indirect cost = 1876
=2447-571
Correct answer is option a. > 1876
Answer 2
Direct labour $13
Direct material 9
Applied manufacturing overhead 8
Unit production (budget) 31184
Unit production (actual) 23971
Actual overhead costs incurred $132,674
Applied overhead = 191768
=23971*8
Actual overhead cost = $132,674
Over applied - overhead $59,094
=191768-132674
Correct answer is option d. Overapplied by 59094